OpenStudy (anonymous):

Where does cscx=1/2 and csc=2?

OpenStudy (anonymous):

unlikely that \(\csc(x)=\frac{1}{2}\) since that would make \(\sin(x)=2\) which would be a miracle

OpenStudy (anonymous):

\[csc^{-1} 1/2 = \frac{1}{\sin^{-1} 1/2} = undefined\]

OpenStudy (anonymous):

Okay, thanks! (: